### v1

# Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation

Abstract
Semi-supervised semantic segmentation, which efficiently addresses the limitation of acquiring dense annotations, is essential for 3D scene understanding. Most methods leverage the teacher model to generate pseudo labels, and then guide the learning of the student model on unlabeled scenes. However, they focus only on points with pseudo labels while directly overlooking points without pseudo labels, namely intra-scene inconsistency, leading to semantic ambiguity. Moreover, inter-scene correlation between labeled and unlabeled scenes contribute to transferring rich annotation information, yet this has not been explored for the semi-supervised tasks. To address these two problems, we propose to explore scene coherence for semi-supervised 3D semantic segmentation, dubbed CoScene. Inspired by the unstructured and unordered nature of the point clouds, our CoScene adopts the straightforward point erasure strategy to ensure the intra-scene consistency. Moreover, patch-based data augmentation is proposed to enhance the inter-scene information transfer between labeled and unlabeled scenes at both scene and instance levels. Extensive experimental results on SemanticKITTI and nuScenes show that our approach outperforms existing methods.
Introduction
LiDAR (Light Detection and Ranging) point cloud semantic segmentation, aiming to recognize the class label for each point of an input point cloud scene, is a core task in autonomous driving. Due to the advancement of deep neural networks, tremendous progress has been achieved (Hu et al. 2020; Zhu et al. 2021; Su et al. 2023). However, the excellent performance of most studies depends heavily on large-scale datasets with high-quality point-level annotations. Such dense and delicate point-level labeling is labor-intensive and time-consuming, hindering the deployment of point cloud semantic segmentation models in real-world applications. Therefore, there is a compelling need to reduce the dependence on huge exhaustively labeled datasets.
Figure 1: (a) Generated pseudo-label scenes from unlabeled scenes containing points with and without pseudo labels simultaneously (green and gray). (b) Regardless of whether point-based or voxel-based segmentation networks are employed, resulting in inconsistencies between forward and backward propagation. (c) Annotation information transfer between labeled and unlabeled scenes.
Semi-supervised learning (SSL), a significant machine learning paradigm for alleviating data annotation demands, seeks to utilize a small number of labeled samples alongside a vast collection of unlabeled samples to optimize deep models. Recently, several semi-supervised 3D semantic segmentation methods (Kong et al. 2023b; Jiang et al. 2021; Li, Shum, and Breckon 2023) have been proposed, with a core focus on effectively transferring diverse semantic information. These methods typically first obtain pseudo-labeled scenes from unlabeled data using the teacher network, and then feed the entire pseudo-labeled scenes into the student network for training. Although these methods have reported promising results, two main issues still need to be addressed.
As shown in Figure 1 (a), the pseudo-labeled scene consists of points with and without pseudo-labels simultaneously. During training, all points are involved in the model’s forward propagation, while only points with pseudo-labels are used for backward propagation, and other points are overlooked. Inconsistency between forward and backward propagation within the scene leads to semantic ambiguity, meaning the segmentation model is unable to determine the semantic labels for points without pseudo-labels. Consequently, this limits the effectiveness of semantic information transfer. Both point-based (Hu et al. 2020) and voxel-based (Choy, Gwak, and Savarese 2019; Zhu et al. 2021) methods suffer from this issue, as illustrated in Figure 1 (b).
Diverse semantic information is crucial for the success of semi-supervised learning of the segmentation models (Kong et al. 2023b; Xiao et al. 2022). Thus, many studies focus on developing data augmentation methods (Sheshappanavar, Singh, and Kambhamettu 2021; Zhang et al. 2022) to expand the data distribution. However, these methods typically model the correlation between two scenes, while neglecting the correlation among multiple scenes. As shown in Figure 1 (c), the labeled and unlabeled scenes may exhibit similarities in factors such as spatial structures. Exploring the correlations between scenes can be beneficial for annotation information transfer, thereby enhancing learning on unlabeled scenes.
To address the above two issues, we explore Scene Coherence and propose CoScene for semi-supervised 3D semantic segmentation. Specifically, based on the unstructured and unordered nature of point clouds, we design a point erasure strategy that removes the points without pseudo labels during forward propagation, achieving full annotation of the training scene and ensuring intra-scene consistency. This strategy complements the insufficiency of the basic pseudo-label mechanism, which can serve as a plug-and-play module for semi-supervised 3D segmentation frameworks based on pseudo labels. Additionally, we design a patch-based data augmentation to model the correlation between multiple scenes at both the scene and instance levels. The key idea is to construct two patch pools to store scene-level and instance-level information, and then combine scene-level and instance-level patches from multiple scenes to generate diverse training samples, facilitating the transfer of annotation information to unlabeled scenes.
The proposed CoScene has been extensively evaluated through experiments and has demonstrated superiority across various benchmarks for semi-supervised 3D semantic segmentation. Under the setting of 5% labeled scenes, our method outperforms the state-of-the-art method by 4% in terms of mIoU on the SemanticKITTI (Behley et al. 2019) dataset. To summarize, our contributions are as follows:
  • We propose a plug-and-play point erasure strategy complementing with pseudo-label mechanism in semi-supervised point cloud semantic segmentation, ensuring intra-scene consistency.
  • We propose a patch-based data augmentation to model inter-scene correlation information at scene and instance levels simultaneously, facilitating information transfer between labeled scenes and unlabeled scene.
  • Extensive experiments demonstrate that our semi-supervised 3D segmentation methods achieves state-of-the-art performance, and comprehensive ablations studies are conducted to demonstrate its effectiveness.
Figure 2: Overview of the proposed CoScene training pipeline. The teacher network generates the pseudo-labels $y_{i}^{u}$ from the unlabeled scene $x_{i}^{u}$, followed by an erasure operation to filter out low-confidence points, ensuring training consistency. For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ a patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, separately. Subsequently, both scenes are fed into the student network for computing the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
Related Work
LiDAR Semantic Segmentation
LiDAR semantic segmentation aims to recognize the class label for each point of an input point cloud scene. Through the development of various deep neural networks, numerous approaches for LiDAR semantic segmentation have been explored over the past few years, which can be broadly grouped into three categories, i.e., voxel-based (Zhu et al. 2021), 2D projection-based (Kong et al. 2023a; Zhao, Bai, and Huang 2021), and point-based methods (Hu et al. 2020; Qi et al. 2017). Prior fully supervised works have made significant progress and achieved appealing performance, while the results deeply depend on large-scale precisely annotated datasets (Behley et al. 2019; Fong et al. 2022). To reduce annotations of 3D data, some methods with less supervision have been explored recently, e.g., weakly-supervised (Pan et al. 2024; Liu et al. 2022; Unal, Dai, and Van Gool 2022; Zhang et al. 2021; Chen et al. 2023; Tang et al. 2024) and semi-supervised learning (Li, Shum, and Breckon 2023; Reichardt, Ebert, and Wasenmüller 2023; Zou et al. 2018; Chen et al. 2021). LaserMix (Kong et al. 2023b), as the pioneer in semi-supervised LiDAR segmentation leverages the simple laser beam mix strategy to get solid results. Follow-up work fuses LiDAR-Camera data to mine deeper information (Chen et al. 2024). For semi-supervised approaches, the overall architecture often builds upon the mean-teacher paradigm (Tarvainen and Valpola 2017) and leverages the concept of pseudo label (Arazo et al. 2020). However, many points without pseudo labels contribute equally to pseudo-label points in the forward propagation, but are directly ignored when performing backward propagation, which brings intra-scene inconsistency issues. Hence, our CoScene leverages the point erasure strategy to mitigate this inconsistency problem, creating a complement to the pseudo-label mechanism to enable effective information transfer.
3D Point Cloud Data Augmentation
Point cloud data augmentation has been explored across various 3D vision tasks (Zhu, Fan, and Weng 2023; Shin et al. 2023; Wang et al. 2024; Fang et al. 2021). Enlightened by the success of Mixup (Zhang et al. 2017) in image classification, several works explore the techniques for mixing selected point cloud samples. PointMixup (Chen et al. 2020) leverages interpolation to generate mixed samples for training. To preserve structural information, RSMix (Lee et al. 2021) provides a neighboring function, and PointCutMix (Zhang et al. 2022) establishes the correspondence relationship for sample fusion. Notably, these works aim at local object-level augmentation. As a result, several works also study the mixing strategies for point cloud scene augmentation. Mix3D (Nekrasov et al. 2021) concatenates points of two scenes to create a new training sample. PloarMix (Xiao et al. 2022) swaps points within the fan-shaped area between two scenes. CoSMix (Saltori et al. 2023) aims to mitigate the domain shift by mixing semantic point groups. These works only support similarity modeling between two scenes, but fail to construct correlation among multiple scenes. By contrast, our approach has the capability to model inter-scene information from multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enhance scenes. Therefore, our approach can effectively guide the transfer of annotation information from labeled scenes to unlabeled scenes, facilitating semi-supervised learning.
Method
In this section, we first state the definition of semi-supervised 3D semantic segmentation and introduce the mainstream teacher-student framework. Then, we present the intra-scene consistency and inter-scene correlation of CoScene, along with its core component, i.e., the point erasure strategy and patch-based data augmentation. Finally, we outline the overall training objective.
Preliminary
Problem Formulation. Before we describe our model and training process, we first provide the definition of semi-supervised point cloud segmentation. In detail, we have a small labeled training set $D_{l}=\left\{x_{i}^{l},y_{i}^{l}\right\}_{i=1}^{N_{l}}$ and a large unlabeled training set $D_{u}=\left\{x_{i}^{u}\right\}_{i=1}^{N_{u}}$, where $x_{i}\in\mathbb{R}^{m\times{(3+r)}}$ is a point cloud scene of $m$ points with three-dimensional Cartesian coordinates relative to the sensor and additional $r$-dimensional feature (e.g., intensity, color, or reflectivity), $y_{i}$ represents its corresponding point-level segmentation labels, $N_{l}$ and $N_{u}$ are the numbers of labeled and unlabeled point cloud scenes, respectively, and $N_{u}\gg N_{l}$ in ideal cases. The goal of SSL-based point cloud segmentation is to learn a segmentation model using the limited labeled data and large-scale unlabeled data.
Teacher-Student Paradigm. Following current mainstream research (Li, Shum, and Breckon 2023), our learning framework also builds upon the teacher-student paradigm, which includes two 3D segmentation networks with the same architecture. Here, we adopt Cylinder3D (Zhu et al. 2021) as the segmentation backbone $\Phi(\cdot)$ to enable a fair comparison with other methods. In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) (Tarvainen and Valpola 2017) of the student weights:
$\theta_{\mathbb{t}}^{k+1}=\alpha\cdot\theta_{\mathbb{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbb{s}}^{k},$ (1)
where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbb{t}}$ and $\theta_{\mathbb{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
Point Erasure Strategy
Based on the teacher-student mutual learning framework (Sohn et al. 2020), the overall architecture of CoScene is depicted in Figure 2. Specifically, CoScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes. How to effectively leverage a large number of unlabeled scenes is crucial for semi-supervised learning. For the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ (Kong et al. 2023b). As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision. This leads to intra-scene inconsistency, which limits the effectiveness of semantic information transfer. To address this issue, we propose erasing points without pseudo labels, resulting in an erased scene $(\hat{x}_{i}^{u},\hat{y}_{i}^{u})$ that contains fully annotated information, similar to fully supervised scenes. Formally, intra-scene inconsistency addressed by the point erasure strategy can be defined as:
$\hat{x}_{i}^{u}=\{x_{i}^{u}|\Phi_{s}(x_{i}^{u})\geq\tau_{s}\},$ (2)
where $\hat{x}_{i}^{u}$ is the erased scene with corresponding class labels $\hat{y}_{i}^{u}$, $\Phi_{s}(\cdot)$ represents the segmentation network that predicts the confidence score for each point in the scene $x_{i}^{u}$, and $\tau_{s}$ is a predefined confidence threshold, equivalent to the pseudo-label mining threshold, requiring no additional adjustment. Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism (see Figure 5 (a)).
Figure 3: Illustration of the pool initialization process. For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enable the extraction of instances belonging to different classes. Subsequently, these patches and instances are stored individually in the patch pool and instance pool.
Patch-based Data Augmentation
Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the transfer of annotation information, guided by inter-scene correlation. The key idea is to construct patch pools to store scene-level and instance-level information, and then combine patches from multiple scenes to generate diverse training samples.
Pool Initialization. We first use the labeled training set $D_{l}$ to create two patch pools, which store local-scene and single-instance information, respectively, as shown in Figure 3. Given the $i$-th point cloud scene $x_{i}^{l}$, we patchify $x_{i}^{l}$ into a sequence of $T$ regular, non-overlapping point patches $\{p_{ij}^{l}\}_{j=1}^{T}$, each of size $(n\times n)$, in the bird’s eye view. This process can be expressed as:
$\left\{p_{ij}^{l}\right\}_{j=1}^{T}=\text{patchify}(x_{i}^{l}),$ (3)
where $p_{ij}^{l}$ denotes the $j$-th point patch from the sequence. From all labeled scenes, we then construct the patch pool $\mathcal{B}_{\text{Pat}}^{l}$, which can be defined as:
$\mathcal{B}_{\text{Pat}}^{l}=\left\{\left\{p_{ij}^{l}\right\}_{j=1}^{T}\right\}_{i=1}^{N_{l}}.$ (4)
Based on the patch pool, we can obtain the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ by assigning different class labels to each patch, which contains instance information for different categories (e.g., car and person). Following the same steps used for labeled scenes, we further create a pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$ and a pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ from the erased scene $\hat{x}_{i}^{u}$. Unlike $\mathcal{B}_{\text{Pat}}^{l}$ and $\mathcal{B}_{\text{Ins}}^{l}$ covering the whole training set $D_{l}$, the pseudo pools $\mathcal{B}_{\text{Pat}}^{u}$ and $\mathcal{B}_{\text{Ins}}^{u}$ only contain scene and instance information from the current batch and are reset in each iteration. This prevents low-quality pseudo labels from persisting in pools and misleading model training. Additionally, if the number of points within a patch or instance falls below a certain threshold $\tau_{\text{min}}$, it will not be saved into the pool, ensuring effective inter-scene information transfer.
Figure 4: (a) Illustration of the MixPatch process. (b) Illustration of selecting instances from the patch pool for filling. An instance will not be filled if there are no other points around the filling location or if the filled instance overlaps with existing instances.
MixPatch. To model the inter-scene correlation and then guide the transfer of annotation information between labeled scenes and unlabeled scenes, we combine multiple scenes using patch pools, as shown in Figure 4. Figure 4 (a) illustrates the patch mixing (MixPatch), while Figure 4 (b) presents the instance filling (InsFill), which is introduced in the next paragraph. During a training iteration, the mini-batch is comprised of labeled and unlabeled scenes. After the unlabeled scene $x_{i}^{u}$ is transformed into the erased scene $\hat{x}_{i}^{u}$ using the point erasure strategy, we integrate information from the patch pool into $\hat{x}_{i}^{u}$. The patch mixing process can be described as follows:
$\hat{x}_{i}^{up}=(\hat{x}_{i}^{u}\odot\mathcal{M})\oplus(\mathcal{B}_{\text{Pat}}^{l}\odot(1-\mathcal{M})),$ (5)
where $\mathcal{M}$ is a binary mask indicating which patches from $\hat{x}_{i}^{u}$ will be replaced with selected patches of the same patch index from the patch pool, $\odot$ denotes element-wise multiplication, and $\hat{x}_{i}^{up}$ represents the augmented scene. For labeled scenes, MixPatch is applied based on the current pseudo patch pool $\mathcal{B}_{\text{Pat}}^{u}$, which can be expressed as the following:
$x_{i}^{lp}=(x_{i}^{l}\odot(1-\mathcal{M}))\oplus(\mathcal{B}_{\text{Pat}}^{u}\odot\mathcal{M}),$ (6)
where $\mathcal{M}$ is the same as defined in Eq. (5), thereby ensuring the full utilization of scene information from the unlabeled scenes. The binary mask $\mathcal{M}$ and its complement $1-\mathcal{M}$ are used to combine inter-scene point cloud patches with the erased scene, facilitating the transfer of annotation information between labeled and unlabeled scenes. To alleviate bias in the patch mixing process, we apply a uniform sampling strategy on the patches of each scene, ensuring that each patch has an equal opportunity to be selected in the mixed scene.
InsFill. In addition to patch mixing, we utilize the instances stored in the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ to further enhance the scene $\hat{x}_{i}^{up}$. Given a patch with location index $j$ in the scene, we randomly select an instance from the instance pool $\mathcal{B}_{\text{Ins}}^{l}$ that was extracted from the $j$-th patch. Patches with identical indices typically correspond to similar LiDAR spatial positions, ensuring consistency between the selected instance and its fill positions in the scene. As depicted in Figure 4 (b), if the selected instance overlaps with existing instances in the scene or lacks sufficient contextual information, the filling operation is bypassed. This judgment process can be easily implemented using the maximum 3D bounding boxes. This data augmentation provides single-instance semantic information for the point cloud scene, enabling the augmented scene $\hat{x}_{i}^{upf}$ and corresponding labels $\hat{y}_{i}^{upf}$ to effectively support semi-supervised training. For the labeled branch, we also apply the same filling strategy as described, selecting instances from the pseudo instance pool $\mathcal{B}_{\text{Ins}}^{u}$ to fill the current scene. This process yields the scene $x_{i}^{lpf}$ along with its corresponding label $y_{i}^{lpf}$.
Model Training
Following the standard teacher-student mutual learning paradigm (Kong et al. 2023b; Sohn et al. 2020), we construct a mini-batch of data with an equal composition of labeled and unlabeled scenes at each iteration. The unlabeled scenes are assigned with pseudo-labels by the teacher model and then train the student model jointly with labeled scenes. We use three types of loss functions to supervise the model training here, which is shown as follows:
$\ \mathcal{L}_{s}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{l}),y_{i}^{l}),$ (7)
$\ \mathcal{L}_{l}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(x_{i}^{lpf}),y_{i}^{lpf}),$ (8)
$\ \mathcal{L}_{u}=\mathcal{L}_{seg}(\Phi_{\theta_{s}}(\hat{x}_{i}^{upf}),\hat{y}_{i}^{upf}),$ (9)
where $\mathcal{L}_{s}$, $\mathcal{L}_{l}$, and $\mathcal{L}_{u}$ denote the segmentation loss for the labeled scene $x_{i}^{l}$, augmented labeled scene $x_{i}^{lpf}$, and augmented unlabeled scene $\hat{x}_{i}^{upf}$, respectively, while $\mathcal{L}_{seg}$ is the original loss of the segmentation network $\Phi(\cdot)$. We sum up all the losses together as our final training target that
$\mathcal{L}=\mathcal{L}_{s}+\lambda_{u}\mathcal{L}_{u}+\lambda_{l}\mathcal{L}_{l},$ (10)
where $\lambda_{u}$ and $\lambda_{l}$ are loss weights that balance the two loss terms. After that, we update the teacher model weights via the EMA strategy described in Eq. (1).
SemanticKITTI nuScenes
Method Venue 1% 5% 10% 20% 40% 50% 1% 10% 20% 50%
Cylinder3D CVPR 2021 45.4 45.4 56.1 57.8 58.7 58.7 50.9 65.9 66.6 71.2
MeanTeacher NeurIPS 2017 45.4 - 57.1 59.2 - 60.0 51.6 66.0 67.1 71.7
CBST ECCV 2018 48.8 - 58.3 59.4 - 59.7 53.0 66.5 69.6 71.6
CPS CVPR 2021 46.7 - 58.7 59.6 - 60.5 52.9 66.3 70.0 72.5
GPC ICCV 2021 - 41.8 49.9 58.8 59.9 - - - - -
ScribbleKITTI CVPR 2022 - 49.9 58.7 59.1 60.9 - - - - -
LiM3D CVPR 2023 58.4 59.5 62.2 63.1 63.3 63.6 - - - -
CylinderMix* AAAI 2024 51.3 56.5 60.8 62.0 62.7 62.5 55.7 69.5 72.0 73.3
LaserMix CVPR 2023 50.6 56.7 60.0 61.9 62.3 62.3 55.3 69.9 71.8 73.2
Ours (CoScene) - 54.5 62.1 63.3 63.7 63.8 64.3 56.6 70.2 72.8 73.9
Table 1: Experimental results on the SemanticKITTI and nuScenes dataset compared with different SSL methods. The results are reported with mIoU scores (%) and the best score for different split is marked in bold. “*” denotes that we reproduce the CylinderMix in the LiDAR branch (Chen et al. 2024) for fair comparison.
Experiments
Datasets and Evaluation Metrics
SemanticKITTI. Following state-of-the-art methods (Kong et al. 2023b; Chen et al. 2024), we evaluate our CoScene on the SemanticKITTI dataset (Behley et al. 2019), which is a large-scale segmentation dataset consisting of LiDAR data derived from the popular KITTI dataset (Geiger, Lenz, and Urtasun 2012). It is composed of 22 sequences, with sequences 00-10 serving as the train set (of which 08 as the val set) and sequences 11-21 as the test set. Moreover, all the point cloud scenes are collected from real-world rural and urban road environments in Germany, and the semantic class labels are mapped to 19 classes in our experiments.
nuScenes. The nuScenes dataset (Fong et al. 2022) is another well-known large autonomous driving dataset. This dataset is composed of 1000 sequences of 20s duration (850 for training and validation, 150 for testing), which are captured by the LiDAR sensor from Boston and Singapore. Following the official splitting protocol, the train and val sets contains 28,130 and 6,019 point cloud scenes, respectively, including per-points annotation for 16 semantic classes.
Evaluation Metrics. For both the SemanticKITTI and nuScenes datasets, we sample different ratios of annotated train scenes to obtain semi-supervised data and ensure the sampling indices are the same as in LaserMix(Kong et al. 2023b) for fair comparison. Meanwhile, we report the performance on the validation set using the mean Intersection over Union (mIoU) metric as the previous works (Li, Shum, and Breckon 2023; Zhu et al. 2021).
Implementation Details
We conduct experiments with 4$\times$ NVIDIA V100 GPUs with 32GB RAM, based on MMDetection3D (Contributors 2020). Inspired by the GT-sampling strategy (Yan, Mao, and Li 2018), we set $\tau_{min}$ to 5 in our experiments. For the threshold $\tau_{s}$ for pseudo-label generation and point erasure, we use the same parameter setting as in LaserMix. To facilitate patch-based data augmentation, we restrict our focus to point clouds within the X and Y dimensions in the range of $[-50,50]m$ for both the SemanticKITTI and nuScenes datasets. Similar to mainstream work, we also adopt the consistency loss proposed in MeanTeacher (Tarvainen and Valpola 2017) to align the predictions of the teacher and student networks. Additionally, the weights of the consistency loss are assigned as $1\times 10^{3}$ and $2\times 10^{3}$ for the SemanticKITTI and nuScenes datasets, respectively. The weights for $\lambda_{u}$ and $\lambda_{l}$ are set as 1 for both datasets.
Main Results
In this subsection, we compare our method with current SOTA semi-supervised point cloud segmentation methods, using sub-sampling of the datasets under various partition protocols (Kong et al. 2023b). Specifically, in the SemanticKITTI dataset (Behley et al. 2019), we split the entire training set into labeled subsets with ratios of 1%, 5%, 10%, 20%, 40%, and 50%. Similarly, in the nuScenes dataset (Fong et al. 2022), we utilize labeled subsets with ratios of 1%, 10%, 20%, and 50%. We provide detailed comparative results on both datasets against different baseline methods.
Results on SemanticKITTI Dataset. We present the comparison results between our method and state-of-the-art methods on the widely adopted SemanticKITTI dataset in Table 1. It can be clearly seen that our method effectively exploits unlabeled scenes compared to the fully supervised learning baseline Cylinder3D (Zhu et al. 2021), with a dramatic performance boost. Specifically, in the 5% ratio, our approach surpasses the fully supervised baseline by 16.7%. In other settings, our method also shows consistent improvements ranging from 5.6% to 7.2% for ratios of 10%, 20%, and 50%. In addition, comparison to the recently proposed SOTA method LaserMix (Kong et al. 2023b), which is considered a strong SSL-based point cloud segmentation method, our approach achieves improvements of 2% to 6% in all cases. The results also demonstrate that our approach outperforms the current SOTA LiM3D (Li, Shum, and Breckon 2023) in the majority of settings. In some partition protocols, our approach shows superior performance compared to the LiM3D with fewer labeled scenes. For example, our approach trained with 10% labeled scenes outperforms LiM3D trained with 40% labeled scenes. It is worth mentioning that our approach utilizes uniform sampling, similar to LaserMix, to acquire labeled scenes, without utilizing the keyframe sampling method used in LiM3D (Li, Shum, and Breckon 2023), which typically yields higher gains when the amount of labeled data is relatively low.
Results on nuScenes Dataset. As shown in Table 1, our approach still shows a significant improvement in performance compared to state-of-the-art methods on the comprehensive nuScenes dataset (Fong et al. 2022). For instance, our approach outperforms LaserMix (Kong et al. 2023b) by 1.3% and 1.0% at 1% and 20% ratios, respectively, demonstrating the generalization ability of our method.
Ablation Study
In this subsection, we study the effects of intra-scene consistency assurance through the point erasure strategy (PtErase), as well as the exploration of inter-scene correlation, incorporating patch-based mix (MixPatch) and patch-based instance filling data augmentation (InsFill) in our method. Without loss of generalizability, all experiments are conducted on the SemanticKITTI dataset with 1%, 10%, 20% and 50% ratio, and evaluated on val set. Table 2 demonstrates the improvements achieved by each of these components.
Ours CoScene Splits on SemanticKITTI Avg.
PtErase MixPatch InsFill 1% 10% 20% 50%
45.4 57.1 59.2 60.0 55.4
51.7 59.8 61.8 60.4 58.4
51.3 61.4 60.7 60.5 58.5
52.3 61.7 60.1 61.1 58.8
54.6 61.0 62.9 62.8 60.3
52.0 61.7 60.8 62.6 59.3
50.5 62.5 63.8 63.6 60.1
54.5 63.3 63.7 64.3 61.5
Table 2: Ablation study of different components of our designed CoScene. PtErase refers to the point erause strategy, while MixPatch and InsFill refers to the patch mix and instance filling data augmentation, respectively. All mIoU results are reported in percentage (%).
Figure 5: Performance comparison on SemanticKITTI val split under different semi-supervised settings. (a) Illustration of performance differences whether using consistency assurance strategy on the LaserMix (Kong et al. 2023b). (b) Comparison between dual-scene level and multi-scene level for the patch-based data augmentation in our CoScene network. (c) Results of mIoU against the number of patch split $n$ under the 10% SemanticKITTI split.
Effectiveness of point erasure strategy. The row 1 in Table 2 denotes the performance of MeanTeacher (Tarvainen and Valpola 2017) as the baseline, revealing a marginal enhancement compared to the supervision-only Cylinder3D model (Zhu et al. 2021). When PtErase is applied to the baseline, it is effective and can boost the performance by ensuring consistency during forward and backward propagation, thereby compensating for the shortcomings of the pseudo-label mechanism. The consistency assurance is implemented through a plug-and-play point erasure strategy, which can be easily integrated into methods based on the pseudo-label mechanism. To further validate the significance of ensuring consistency during forward and backward propagation, we embedded the PtErase strategy into LaserMix. As illustrated in Figure 5 (a), performance improves across all ratios when using the PtErase strategy. This further emphasizes the importance of intra-scene consistency and demonstrates the generalization ability of our method.
Effectiveness of patch-based data augmentation. The patch-based data augmentation consists of MixPatch (patch-based mix) and InsFill (instance filling). Rows 3 and 4 of Table 2 demonstrate that using MixPatch or InsFill alone results in an average mIoU improvement of 3.1% or 3.4%, respectively, compared to the baseline. This suggests that annotation information transfer across scenes benefits SSL 3D segmentation. As shown in rows 5 and 6 of the Table 2, combining MixPatch or InsFill with PtErase significantly enhances performance, which indicates the effectiveness of the patch-based data augmentation. Furthermore, the collaboration of all three components, leads to substantial performance gains, as validated by the last row of Table 2. Additionally, our patch pool and instance pool within the patch-based data augmentation facilitate effective information fusion across multiple scenes. This approach inherently enables the transfer of inter-scene diverse correlation information. To verify the effectiveness of the interaction at the multi-scene level, we modify the pool implementation to enable only interaction between two scenes. The comparative results, as shown in Figure 5 (b), indicate superior performance at multi-scene level compared to dual-scene level.
Different mixing strategies. Figure 5 (c) compares the patch-based data augmentation employed in our CoScene with other mixing approaches (Nekrasov et al. 2021; Xiao et al. 2022; Yun et al. 2019; Zhang et al. 2017). MixUp demonstrates a slight improvement over the baseline, likely due to its limited ability to transform information. Other methods have a considerable improvement over the baseline by enriching point cloud distributions. In contrast, our CoScene achieves the greatest gain by effectively exploring rich inter-scene correlation, highlighting its effectiveness in facilitating information transfer between labeled and unlabeled scenes. This is accomplished through its innovative design, which includes a patch pool and an instance pool spanning multiple scenes.
Patch split number. As the patch split number $n$ plays a critical role in our framework, we further analyze the impact of varying $n$ in patch-based data augmentation, with the results presented in Figure 6. When $n$ is small, it can yield a certain level of enhancement, but it may fail to generate data sufficiently divergent from the original, impeding the effective transfer of information between labeled and unlabeled scenes. Conversely, when $n$ is too large, the generated scenes may deviate excessively from the original data, resulting in the loss of essential prior scene information. Empirically, we identify that setting $n$ to 18 yields optimal results. Therefore, throughout this paper, we consistently use $n$ set to 18, unless explicitly stated otherwise.
Figure 6: Comparison of different mixing-based techniques.
Conclusion
In this paper, we explore the scene coherence for semi-supervised point cloud segmentation, building upon the intra-scene consistency and inter-scene correlation. In particular, we achieve intra-scene consistency assurance through a point erasure strategy that ensures consistency during the forward and backward propagation, rather than simply disregarding points with pseudo-labels as in previous methods. Additionally, we explore inter-scene correlation and propose a patch-based data augmentation to enhance information transfer between labeled scenes and unlabeled scenes. Benefiting from these two components, our CoScene effectively transfers diverse semantic information and produces more accurate segmentation results. Extensive experimental results demonstrate the superiority of our CoScene over state-of-the-art methods.